#6. A group of workers exposed to high levels ofCO were categorized on the ymptomrshin Morning Evenin Influenza Hendache, 12 . weakness-ㄧㄧㄧㄒㄏㄧ˙-1T6 Shortness of Breath Suppose that a worker is chosen at random (o) What is the probability that a morning shift worker suffered headaches (b) What is the probability that a worker suffered weakness or worked nighte (c) For this population of workers are symptoms independent of shift? Explain.
Answer #1

Solution :

Total = 16 + 13 + 18 + 24 + 33 + 6 + 11 + 16 + 5 + 7 + 9 + 9 = 167

(a)

P(Headache) = (24 + 33 + 6) / 167 = 63 / 167 = 0.3772

Probability = 0.3772

(b)

P(weakness or worked nights) = (11 + 16 + 5) / 167 + (18 + 6 + 5 + 9) / 167 - 5 / 167

= 32 / 167 + 38 / 167 - 5 / 167

= (32 + 38 - 5) / 167

= 65 / 167

= 0.3892

Probability = 0.3892

(c) Yes,